OpenStudy (anonymous):

I give best response if you answer these 5 questions!!!!!

OpenStudy (anonymous):

Write a ratio in simplified form of the vertical length to the horizontal length for the red triangle

OpenStudy (anonymous):

Please!! Help!!!

OpenStudy (anonymous):

vertical length is 2, horizontal length is 3, ratio is \(\frac{2}{3}\)

OpenStudy (anonymous):

The red triangle goes 2 units up vertically for 3 units it goes to the right horizontally. Thus the ratio is: 2/3.

OpenStudy (anonymous):

Okay Would a triangle with a vertical length of 6 and a horizontal length of 10 have the same slope as the blue and red triangles shown in the graph?

OpenStudy (anonymous):

6/10 = 3/5 --> This is not the same as 2/3. Hence the slopes are different.

OpenStudy (anonymous):

so what would the slope be of 6/10?

OpenStudy (anonymous):

6/10 is a fraction which can be reduced to 3/5. This means that the line goes 3 units up for every 5 units it goes to the right, so slope is 3/5. Btw, try solving some of these yourself? You will never learn otherwise...

OpenStudy (anonymous):

You can ask for help when you really need it of course.

OpenStudy (anonymous):

if u want your Best Responce id not be telling me that cause im not forced to its a want to thing!

OpenStudy (anonymous):

How does the slope of the red triangle compare to the slope of the blue triangle?

OpenStudy (anonymous):

sorry here is the next question

OpenStudy (anonymous):

Which equation would you use to find out if the two lines in the graph are parallel?